Solve the equation by factoring. (Enter your answers as a comma-separated list.)


2x2 + 11x − 21 = 0

Answer by Alan3354(69443)


2x^2 + 11x - 21 = 0
============
It'll be (ax + b)*(cx + d)
a*c = 2
b*d = -21
-------------
Factoring is trial and error.
Nothing is gained by someone doing it for you.